OpenStudy (anonymous):

Can someone help me with two problems What is the solution to the system of equations? y = 2x + 6 3x - y = 6

Directrix (directrix):

y = 2x + 6 Equation 1 3x - y = 6 Equation 2 ======== Take equation 1 and for y in equation 2, substitute for y.

Directrix (directrix):

3x - y = 6 Replace y by (2x + 6) 3x - (2x + 6) = 6 ----> Solve this for x and post what you get, okay? @Soniatonia

OpenStudy (anonymous):

kk

Directrix (directrix):

Did you get x?

OpenStudy (anonymous):

x - 6 = 6

Directrix (directrix):

No. 3x - (2x + 6) = 6 3x - 2x - 6 = 6 What next? @Soniatonia

OpenStudy (anonymous):

I think you simplify

Directrix (directrix):

Yes, and what do you get?

Directrix (directrix):

What is 3x - 2x =

OpenStudy (anonymous):

x

Directrix (directrix):

3x - 2x - 6 = 6 x - 6 = 6 x = 6 + 6 x = 12

Directrix (directrix):

Your task is to find y.

OpenStudy (anonymous):

k

Directrix (directrix):

In this equation y = 2x + 6, in place of x, substitute x = 12 and solve for y.

Directrix (directrix):

Post what you get.

OpenStudy (anonymous):

y = 2 (x + 3)

Directrix (directrix):

y = 2 * 12 + 6

Directrix (directrix):

So, what does that equal?

Directrix (directrix):

Note: In this >> y = 2 (x + 3), you gave the factorization of 2x + 6. That's not the process to solve for x here. But, it's good that you factored correctly.

OpenStudy (anonymous):

y=30

Directrix (directrix):

@Soniatonia Two times Twelve and then add Six and you will have y.

Directrix (directrix):

Correct. y =30 So, the solution to the pair of equations is written like this: ( 12, 30 )

OpenStudy (anonymous):

ty you so much

Directrix (directrix):

You are welcome;